OpenStudy (anonymous):

-80+5b= -4(b-7) solve

OpenStudy (anonymous):

-80+5b=-4(b-7) -80+5b=-4b+28 +4b +4b -80+9b=28 +80 +80 9b=108 /9 /9 b=12